Recent Log Entries

Pathlength Correction up to 7R

Log Number: 4230141Logbook(s): Entry Makers: moserBody: 

C-beam to 7R dumplette. Figure 1: Before. Figure 2: DogCalc. Figure 3: Not downloading Dog1 so we don't have to re-check gangphases, dialed in shortly before this. Figure 4: After corrections. Good enough for now.

RATN Trim Up/Down

Log Number: 4230136Logbook(s): Tags: Body: Ran from opsl05

Finished after 5 iterations.

GREY: RF OFF
RATNI RED: <= 3.0 dB
GASKL RED: <1.8 V or > 2.4 V

LEM North Linac

Log Number: 4230135Logbook(s): Tags: Body: LEM 9.17 was successfully run for the North Linac on 2023-12-05 05:59:37. The application phase took 00:01:13 to complete.

LEM South Linac

Log Number: 4230134Logbook(s): Tags: Body: LEM 9.17 was successfully run for the South Linac on 2023-12-05 05:58:47. The application phase took 00:00:15 to complete.

Phaser Results

Log Number: 4230133Logbook(s): Body: Results where Job # is 801 {9} CAVITY PHASE ERROR (DEGREES) OUTCOME PHASE (DEGREES) START DATE DURATION (SECONDS) CORRECTION DATE CORRECTION ERROR REASON 2L22-1 -5.36 CORRECTED 38.20 05-Dec-2023 05:39 56</t

Cold RF ON 1L25-2

Log Number: 4230132Logbook(s): Entry Makers: bjacobBody: 

This cavity gave us a few trips so we figured a POFF optmization was in order. I decided to try Cold RF ON and see if that was able to quickly fix our problem. It only made a minor adjustment to POFF but we'll keep an eye on it and see if the cavity is looking better.

Threading Beam to 2S

Log Number: 4230130Logbook(s): Entry Makers: moserBody: 

Figure 1: Initial beam into SL not quite making it to 2L10 BPM.

2L05 was recovered since last beam through SL, so starting with it's zone gang phase first. Figure 2. Then, moving on to SLGP.

Moved about 100deg of SLGP, slowly gaining more and more BPMs, all the way to 2L13. But then tripped 2L13 BLVF. Typing this up as we wait for vacuum to recover.

All adjustments done in Figure 2 were w/VL-mode from C-laser.

Figure 3: Getting towards the end of the linac. Figure 4: At the end of the SL, now trying to maximize 2S BPM wiresums by adjust SLGP. Figure 5: Current orbit through 2L.

1L22 viewer not working

Log Number: 4230126Logbook(s): Entry Makers: yoonBody: 

viewer isnt working, tried reinserting a few times with no success

Late Entry: MBZ5E02 Shunt Mismatch

Log Number: 4230122Logbook(s): Entry Makers: bjacobBody: 

Dog5E shunt dropped to zero current. A small jiggle of the field set point was enough to wake it back up.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 4230118Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: moser
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09<

Pages